Skip to content
CalcGospel 國際數學圖譜
返回

S1 2022 Oct Q6

A Level / Edexcel / S1

IAL 2022 Oct Paper · Question 6

题目

Problem

The Venn diagram shows the events AA, BB, CC and DD, where pp, qq, rr and ss are probabilities.

(a) Write down the value of

(i) P(A)P(A)

(ii) P(A∣B)P(A\mid B)

(iii) P(A∣C)P(A\mid C)

(3)

Given that P(B′∩D′)=710P(B'\cap D')=\dfrac{7}{10} and P(C∣D)=35P(C\mid D)=\dfrac{3}{5}

(b) find the exact value of qq and the exact value of rr

(6)

Given also that P(B∪C′)=58P(B\cup C')=\dfrac{5}{8}

(c) find the exact value of ss

(2)
题目中文翻译

维恩图表示事件 AA、BB、CC 和 DD,其中 pp、qq、rr 和 ss 均为概率。

(a) 写出:(i) P(A)P(A);(ii) P(A∣B)P(A\mid B);(iii) P(A∣C)P(A\mid C)。

已知 P(B′∩D′)=710P(B'\cap D')=\dfrac{7}{10} 且 P(C∣D)=35P(C\mid D)=\dfrac{3}{5}。

(b) 求 qq 和 rr 的准确值。

又已知 P(B∪C′)=58P(B\cup C')=\dfrac{5}{8}。

(c) 求 ss 的准确值。

解答

(a)

解法一

思路

展开

这题主要读 Venn 图。BB 完全在 AA 中,所以 P(A∣B)=1P(A\mid B)=1;AA 与 CC 不相交,所以 P(A∣C)=0P(A\mid C)=0。

答题过程

展开

From the Venn diagram,

P(A)=0.13+0.12=0.25.\begin{align*} P(A)=0.13+0.12=0.25. \end{align*}

Since BB is inside AA,

P(A∣B)=1.\begin{align*} P(A\mid B)=1. \end{align*}

Since AA and CC do not overlap,

P(A∣C)=0.\begin{align*} P(A\mid C)=0. \end{align*}

(b)

解法一

思路

展开

由 P(C∣D)=35P(C\mid D)=\dfrac35 得 qq+r=35\dfrac{q}{q+r}=\dfrac35。再用 P(B′∩D′)=710P(B'\cap D')=\dfrac{7}{10} 和总概率为 11,可推出 q+r=0.18q+r=0.18,然后联立。

答题过程

展开

Using P(C∣D)=35P(C\mid D)=\dfrac{3}{5},

qq+r=35.\begin{align*} \frac{q}{q+r}=\frac{3}{5}. \end{align*}

So

5q=3q+3r,\begin{align*} 5q=3q+3r, \end{align*}

and therefore

q=32r.\begin{align*} q=\frac{3}{2}r. \end{align*}

From the diagram and P(B′∩D′)=710P(B'\cap D')=\dfrac{7}{10},

0.13+p+s=710.\begin{align*} 0.13+p+s=\frac{7}{10}. \end{align*}

Using the total probability,

p+q+r+s+0.12+0.13=1.\begin{align*} p+q+r+s+0.12+0.13=1. \end{align*}

Subtracting the first equation from the total gives

q+r+0.12=0.30.\begin{align*} q+r+0.12=0.30. \end{align*}

So

q+r=0.18.\begin{align*} q+r=0.18. \end{align*}

Now solve with q=32rq=\dfrac32r:

32r+r=0.18.\begin{align*} \frac{3}{2}r+r=0.18. \end{align*}

Hence

r=0.072=9125,\begin{align*} r=0.072=\frac{9}{125}, \end{align*}

and

q=0.108=27250.\begin{align*} q=0.108=\frac{27}{250}. \end{align*}

(c)

解法一

思路

展开

B∪C′B\cup C' 包含 BB 区域、AA only 区域、DD only 区域和外部 ss,不包含 CC only 的 pp 与 qq。

答题过程

展开

From the diagram,

P(B∪C′)=0.13+0.12+r+s.\begin{align*} P(B\cup C')=0.13+0.12+r+s. \end{align*}

Using P(B∪C′)=58P(B\cup C')=\dfrac{5}{8} and r=0.072r=0.072,

58= 0.13+0.12+0.072+ss= 0.303.\begin{align*} \frac{5}{8} =&\,0.13+0.12+0.072+s\\ s=&\,0.303. \end{align*}